### v1

# An iterative construction of complete Kähler–Einstein metrics

Abstract.
We extend Tsuji’s iterative construction of a Kähler–Einstein metric with negative scalar curvature to (non-compact) Kähler manifolds with bounded geometry, using Berndtsson’s method for the compact setting. Consequently, given a holomorphic surjective map $p:X\to Y$, where $X$ is a weakly pseudoconvex Kähler manifold and $Y$ is a complex manifold, whose smooth fibers admit Kähler–Einstein metrics with negative scalar curvature and bounded geometry, we show that the fiberwise Kähler–Einstein metric induces a semi-positively curved metric on the relative canonical bundle $K_{X/Y}$ of $p$. Moreover, our approach can be applied to obtain the plurisubharmonic variation of cusp Kähler–Einstein metrics.
Key words and phrases:
Monge-Ampère operators, Bergman kernels, families of Kähler manifolds
2010 Mathematics Subject Classification:
32U15, 32Q15, 32W20
1. Introduction
In the 1950s, Calabi conjectured that a compact Kähler manifold with negative first Chern class admits a unique Kähler–Einstein metric. This was famously shown independently by Aubin [Aub78] and Yau [Yau78], leading to various remarkable results in algebraic geometry for these manifolds. A question arises as to whether Kähler–Einstein metrics can be the limit of a sequence of algebraic metrics induced by the pluri-canonical line bundle; cf. [Yau87]. Through the Tian–Yau–Zelditch asymptotic expansion of the Bergman kernel, Donaldson [Don01] iteratively constructed on a polarized manifold $(X,L)$ with discrete automorphism group, a sequence of ”balanced metrics” whose limit is a Kähler form with constant scalar curvature. Tsuji [Tsu10] introduced a different approach to obtaining Kähler–Einstein metrics with negative scalar curvature by algebraic approximations.
Beyond the compact setting, it is natural to investigate the existence of complete Kähler–Einstein metrics on non-compact Kähler manifolds. Initially, Cheng–Yau [CY80] established necessary geometric conditions for obtaining complete Kähler–Einstein metrics with negative scalar curvature on non-compact manifolds. Later, this result was extended by Kobayashi [Kob84] and Tian–Yau [TY87]. At the same time, Mok and Yau [MY83] showed the existence of a complete Kähler–Einstein metric with negative scalar curvature on an arbitrary bounded pseudoconvex domain. Recently, Wu and Yau [WY20] imposed an assumption on the holomorphic sectional curvature to construct a unique complete Kähler–Einstein metric with negative Ricci curvature, using a new complex Monge–Ampère equation. We also refer to  [Ton18, HLTT22] for other proofs of Wu–Yau’s result using the Kähler–Ricci flow.
Motivated by Tsuji’s iteration for numerically approximating canonical Kähler metrics in [Tsu10, Tsu13], we study Tsuji’s iterative procedure in the non-compact setting. We aim to obtain uniform convergence for Tsuji’s iterations, making use of the non-compact version of the asymptotic expansion of the Bergman kernels.
To state our result, we first fix some notation and terminology. Let $X$ be a complex manifold of dimension $n$. Let $L$ be a positively curved Hermitian holomorphic line bundle on $X$ with reference curvature form $\omega=dd^{c}\phi_{L}$, where $\phi_{L}$ denotes the weight of a Hermitian metric $h_{L}=e^{-\phi_{L}}$ on $L$. We consider the space $\mathcal{H}_{L}$ of (possibly singular) Hermitian metrics on $L$ with positive curvature current; they can be expressed as $h=e^{-\phi}$ with plurisubharmonic (psh) weight $\phi$, and $\omega_{\phi}$ denotes the associated curvature current. One can also identify a Hermitian metric $h=e^{-\phi}$ on $L$ with a function $\varphi=\phi-\phi_{L}$, so that $h=h_{L}e^{-\varphi}$.
By $H^{0}(X,L)$, we denote the space of globally holomorphic sections of $L$. Let $\mu$ be a positive measure on $X$. For any positive integer $k$, we define the Hermitian product on $H^{0}(X,L^{k})$ with the $L^{2}$ norm as follows:
(1.1) $\|s\|^{2}_{k\phi,\mu}:=\int_{X}|s|^{2}e^{-k\phi}{\rm d}\mu.$
The integrand $|s|^{2}e^{-k\phi}$ can be understood as $|s|^{2}_{h_{L}}e^{-k\varphi}$ with $\varphi=\phi-\phi_{L}$. Then, we define by
(1.2) $B_{k\phi,\mu}(x)=\sup\{|s(x)|^{2}e^{-k\phi(x)}\colon s\in H^{0}(X,L^{k}),\|s\|_{k\phi,\mu}\leq 1\}.$
the Bergman function, and $\mathcal{K}_{k\phi,\mu}=B_{k\phi,\mu}e^{k\phi}$ is the Bergman kernel associated with $(L,k\phi,\mu)$. In particular, $\dfrac{1}{k}\log\mathcal{K}_{k\phi,\mu}$ is defined to be a Hermitian metric on $L$.
Fixing a smooth measure $\Omega$ on $X$, we are interested in the following complex Monge–Ampère equation
(1.3) $\omega^{n}_{\phi_{\infty}}=e^{\phi_{\infty}-\phi_{L}}\Omega,$
where $\phi_{\infty}$ is an unknown smooth Hermitian metric on $L$ with positive curvature. Set $\varphi:=\phi_{\infty}-\phi_{L},\omega:=dd^{c}\phi_{L}$, then the equation (1.3) can be rewritten as
(1.4) $(\omega+dd^{c}\varphi)^{n}=e^{\varphi}\Omega.$
In particular when $\Omega$ satisfies $\operatorname{\textrm{Ric}}(\Omega)=-\omega$, then $\operatorname{\textrm{Ric}}(\omega_{\phi_{\infty}})=-\omega_{\phi_{\infty}}$, hence $\omega_{\phi_{\infty}}$ is a Kähler–Einstein metric with $-1$ Ricci curvature.
We now describe Tsuji’s iteration associated with the equation (1.3). Assume that $\phi$ is a (possibly singular) Hermitian metric on $L$. Let $\mathcal{K}_{k\phi,\mu_{\phi}}$ be the Bergman kernel associated with the weight $k\phi$ and the measure $\mu_{\phi}:=e^{\phi-\phi_{L}}\Omega/n!$. The quantity
(1.5) $\beta_{k}(\phi):=\dfrac{1}{k}(\log\mathcal{K}_{k\phi,\mu_{\phi}}-\log d_{k}),$
defines again a metric on $L$, with
$d_{k}=\left(\frac{k}{2\pi}\right)^{n}.$
Starting from a metric $\phi_{1}$ on $L$, we inductively obtain a sequence of Hermitian metrics on $L$ by setting $\phi_{k+1}:=\beta_{k}(\phi_{k})$. Generalizing the works of Tsuji [Tsu10], Song–Weinkove [SW10] and Berndtsson [Ber09b], we show that $\phi_{k}$ uniformly converges to the metric $\phi_{\infty}$ which satisfies the equation (1.3). In particular, we do not assume the general equation (1.3) is the one for Kähler–Einstein metrics as the previous works.
Theorem 1.1.
Let $X$ be an $n$-dimensional Kähler manifold equipped with a positively curved line bundle $(L,e^{-\phi_{L}})$. Let $\Omega$ be a smooth volume form on $X$. Assume that there exists a smooth Hermitian metric $\phi_{\infty}$ on $L$ satisfying the complex Monge–Ampère equation
(1.6) $\omega^{n}_{\phi_{\infty}}=e^{\phi_{\infty}-\phi_{L}}\Omega.$
Assume moreover that $(X,\omega_{\phi_{\infty}})$ is a Kähler manifold with bounded geometry of order $\ell\geq 5$. Let $\phi_{1}$ be a continuous metric on $L$ such that $\sup_{X}|\phi_{1}-\phi_{\infty}|\leq C$. Let $\phi_{k}$ be a sequence of metrics on $L$, defined by Tsuji’s iteration as described above. Then $\sup_{X}|\phi_{k}-\phi_{\infty}|\rightarrow 0$ as $k\rightarrow\infty$ at the rate $\frac{\log k}{k}$.
Remark 1.2.
In particular, when $L=K_{X}$ is an ample line bundle equipped with a smooth Hermitian metric $e^{-\phi_{L}}$, we can take $\Omega=e^{\phi_{K_{X}}}$ as a smooth volume form. For any (smooth) metric $e^{-\phi}$ on $K_{X}$, the $L^{2}$ norm on $H^{0}(X,K_{X}^{(k+1)})$ is defined by
$\|\sigma\|_{k\phi,\mu_{\phi}}^{2}:=c_{n}\int_{X}\sigma\wedge\bar{\sigma}e^{-(k+1)\phi}e^{\phi}=\int_{X}|\sigma|^{2}e^{-k\phi}.$
Similarly, we define Tsuji’s iteration, which is analogous to that in [Ber09b]. The only difference is that the constant $d_{k}$ is the dimension of $H^{0}(X,K_{X}^{(k+1)})$ in the compact setting. It is known that $\dim H^{0}(X,L^{k})=O(k^{n})$ for ample line bundle $L$, so our definition for $d_{k}$ is the one of Berndtsson modulo a uniform constant.
As mentioned above, one considers the special case where $L=K_{X}$ is ample, it follows from [CY80, WY20] that, under extra conditions, there exists a unique Kähler–Einstein metric $\omega_{\rm KE}$ on $X$ such that $\operatorname{\textrm{Ric}}(\omega_{\rm KE})=-\omega_{\rm KE}$ with
$\omega_{\rm KE}=\operatorname{\textrm{dd}^{c}}\log(e^{u}\omega^{n})=\operatorname{\textrm{dd}^{c}}\log(\omega^{n})+dd^{c}u$
for a bounded smooth function $u$. We denote by $\phi_{\omega}=\log\omega^{n}$ and $\phi_{\rm KE}=\log(e^{u}\omega^{n})$ as two Hermitian metrics on $K_{X}$. Then, we obtain the following Monge–Ampère equation
(1.7) $(\operatorname{\textrm{dd}^{c}}\phi_{\rm KE})^{n}=e^{\phi_{\rm KE}-\phi_{\omega}}\omega^{n}.$
In this case, we define Tsuji’s iteration $\phi_{k+1}=\beta_{k}(\phi_{k})$ with $\beta_{k}$ defined as in (1.5) and
(1.8) $d_{k}=\left(\frac{k}{2\pi}\right)^{n}\max\left\{\frac{1}{2},1+\frac{{\rm Scal}(\omega_{\rm KE})}{2k}\right\}=\left(\frac{k}{2\pi}\right)^{n}\max\left\{\frac{1}{2},1-\frac{n}{2k}\right\}.$
Then, we obtain the non-compact version of Tusji’s iteration (cf. [Tsu10, SW10, Ber09b]).
Theorem 1.3.
With assumptions in Theorem 1.1, suppose further that $L=K_{X}$ with $\phi_{\infty}=\phi_{\rm KE}$ satisfying the equation (1.7). Let $\phi_{1}$ be a continuous metric on $K_{X}$ such that $\sup_{X}|\phi_{1}-\phi_{\rm KE}|\leq C$. Let $\phi_{k}$ be a sequence of metrics on $K_{X}$, defined by Tsuji’s iteration as described above. Then $\sup_{X}|\phi_{k}-\phi_{\rm KE}|\rightarrow 0$ as $k\rightarrow\infty$ at the rate $\frac{1}{k}$.
This result also extends the uniform convergence proved by the second author [Tô22] and, more recently, by Yoo [Yoo24] for pseudoconvex domains with bounded geometry.
The idea of the proof of our theorems is to use the following uniform asymptotic expansion of the Bergman function:
(1.9) $\left|B_{k\phi,\omega_{\phi}^{[n]}}-\left(\frac{k}{2\pi}\right)^{n}\left(1+\frac{b_{1}}{2k}\right)\right|_{\mathcal{C}^{0}}\leq Ck^{n-2}.$
where $\omega^{[n]}:=\frac{\omega^{n}}{n!}$ and $C>0$ only depends on the bounded geometry of $(X,\omega_{\phi})$, extending Berndtsson’s work [Ber09b] in the compact setting. A more general result on the asymptotic expansion (1.9) in $C^{k}$-norms was proved in Ma–Marinescu [MM07, Problem 6.1] (cf. also [MM15, Theorem 6]. For the reader’s convenience, we provide an alternative proof for the uniform Bergman expansion (1.9) (cf. Theorem 3.1) using the method of peak sections introduced by Tian [Tia90] (see also [Rua98, Lu00]) and the Laplace method as in [Don09, Cha15]. We refer to [BdMS76, Tia90, Cat99, Zel98, Lu00, Cha03, DLM06, MM07, BBS08, MM15] and references therein for more details on the asymptotic behavior of Bergman kernels on compact Kähler manifolds.
As a consequence, we apply Theorem 1.3 to study the variation of complete Kähler–Einstein metrics, extending the results in [Tsu10, Sch12, Ber13] to the non-compact setting. We refer to [Pău17, Gue20, CGP21] for more results in the compact case.
Theorem 1.4.
Let $p:X\to Y$ be a holomorphic surjective map from a weekly pseudoconvex Kähler manifold $X$ to a complex manifold $Y$. Let $Y^{\circ}\subset Y$ be a set of points that are not critical values of $p$ in $Y$ and set $X^{\circ}=p^{-1}(Y^{\circ})$ and $X_{y}=p^{-1}(\{y\})$. Assume that
  • (i)
    for every $y\in Y^{\circ}$, $X_{y}$ admits a Kähler–Einstein metric $\omega_{\textrm{KE},y}$ with negative Ricci curvature such that $(X_{y},\omega_{\textrm{KE},y})$ has bounded geometry of order $\ell\geq 5$ which is locally uniform in $y$,
  • (ii)
    $K_{X/Y}$ admits a singular Hermitian metric $h_{0}=e^{-\phi_{0}}$ with semi-positive curvature current and for every $y_{0}\in Y^{0}$, there exists neighborhood $U\Subset Y^{\circ}$ of $y_{0}$ such that $\|\phi_{0,y}-\phi_{{{\rm KE},y}}\|_{L^{\infty}(X_{y})}\leq C(U)$ for every $y\in U$.
Then the Hermitian metric on $(K_{X/Y})|_{X^{\circ}}$ induced by the Kähler–Einstein metrics on the fibers is positively curved. Moreover, if we have the additional condition that
  • (iii)
    the Kähler–Einstein potential $(\phi_{{\rm KE},y})_{y\in Y^{\circ}}$ is bounded from above near the singular fibers,
then this metric extends canonically across $X\setminus X^{\circ}$ to a positively curved metric on $K_{X/Y}$.
As a consequence, we obtain the positive property on the variation of Kähler-Einstein metric on pseudoconvex domain proved in [Tsu13, Cho15b, CY22], following the argument of Tsuji [Tsu13] (cf. Theorem 4.8). See also Choi–Yoo [CY21] for a similar result for strongly pseudoconvex domains contained in a Kähler manifold using the Schumacher method [Sch12].
Furthermore, the statement of Theorem 1.4 also holds in the following setting (cf. Theorem 4.11). Let $p:\overline{X}\to Y$ be a holomorphic surjective proper map between compact Kähler manifolds with relative dimension $n$. Let $X:=\overline{X}\setminus D$ where $D=\sum_{i=1}^{\ell}D_{i}$ is a reduced divisor with simple normal crossing support. Assume $Y^{\circ}$ is the largest Zariski open subset of $Y$ such that if $\overline{X}^{\circ}:=p^{-1}(Y^{\circ})$ then every fiber ${\overline{X}}_{y}:=p^{-1}(\{y\})$ is smooth, $D_{y}:=D\cap{\overline{X}_{y}}$ has simple normal crossing support. We denote by $X^{\circ}=\overline{X}^{\circ}\setminus D$. Then we show that the current obtained by gluing the fiberwise Kähler–Einstein metrics with cusp singularities along $D$ is positive.
Recall that $\omega_{y}$ is said to have cusp singularities (or Poincaré singularities) along $D_{y}$ if, locally $D_{y}=(z_{1}\cdots z_{\ell}=0)$, $\omega_{y}$ is quasi-isometric to the cusp metric:
$\omega_{\rm cusp}:=\sum_{j=1}^{\ell}\frac{idz_{j}\wedge d\bar{z}_{j}}{|z_{j}|^{2}\log^{2}|z_{j}|^{2}}+\sum_{k=\ell+1}^{n}idz_{k}\wedge d\bar{z}_{k}.$
Assume that $K_{\overline{X}_{y}}+D_{y}$ is ample for any $y\in Y^{\circ}$. By the well-known result of Kobayashi [Kob84] and Tian–Yau [TY87], building upon Cheng–Yau’s method [CY80] (see also [Ban90, Wu08, Gue14, BG14, GW16, GP16] for more details), for each fiber, there exists a (unique) complete Kähler metric $\omega_{y}$ on $X_{y}:=\overline{X}_{y}\backslash D_{y}$ with cusp singularities along $D_{y}$ and $\operatorname{\textrm{Ric}}(\omega_{y})=-\omega_{y}$ on $X_{y}$. The metric $\omega_{y}$ can be extended as a positive current on $\overline{X}_{y}$; it satisfies
(1.10) $\operatorname{\textrm{Ric}}(\omega_{y})=-\omega_{y}+[D_{y}]$
in the sense of currents on $\overline{X}_{y}$, where $[D_{y}]$ is the integration current along $D_{y}$.
Over $X^{\circ}$, we can glue the fiberwise Kähler–Einstein metrics $\omega_{y}$ with cusp singularities to obtain a current $T$. Using a similar statement of Theorem 1.4 for this setting (cf. Themrem 4.11), we get:
Theorem 1.5.
With the setup above, the fiberwise cusp Kähler–Einstein metrics $\omega_{y}$ satisfying (1.10) can be glued to define a current $T$ on $\overline{X}$ such that:
  1. (i)
    $T$ is positive on $X^{\circ}=\overline{X}^{\circ}\backslash D$ ;
  2. (ii)
    $T$ extends canonically across $(X\backslash X^{\circ})\cup D$ to a closed positive current on $\overline{X}$.
We remark that a similar result for the family of twisted conic Kähler–Einstein metrics was obtained by Guenancia [Gue20], and a result for the family of cusp Kähler–Einstein metrics was proved by Naumann [Nau22] (see also [Gue20, Section 8.1]). In particular, they used the approach of Schumacher [Sch12], while we followed the method of Tsuji [Tsu10].
Organization of the paper
In Section 2, we start with preliminary materials that we will need concerning (singular) Hermitian metrics on a line bundle and local quasi-coordinates. The main Theorems 1.1 and 1.3 are proved in Section 3, where we also establish an asymptotic expansion for the Bergman kernel associated with weighted spaces. Section 4 is devoted to the study of the positivity variation of complete Kähler–Einstein metrics, where Theorems 1.4 and  1.5 are proved.
Acknowledgement
The authors are grateful to Vincent Guedj and Henri Guenancia for their support and suggestions. We would like to thank Bo Berndtsson for sharing his paper [Ber09b]. We also thank J. Cao, L. Charles, X. Ma, and S. Yoo for their interest in this work and valuable discussions. Part of this work was done while the second author was visiting Vietnam Institute for Advanced Study in Mathematics (VIASM), and he would like to thank VIASM for its hospitality. Q. T. Dang is partially supported by the PARAPLUI ANR-20-CE40-0019 project. T. D. Tô is partially supported by ANR-21-CE40-0011-01 (research project MARGE), PEPS-JCJC-2024 (CRNS) and Tremplins-2024 (Sorbonne University).
2. Notation and background
2.1. General notation
Throughout the paper, $(X,\omega)$ denotes a Kähler manifold of dimension $n$ equipped with a Kähler metric $\omega$. We denote by $d=\partial+\bar{\partial}$ and ${d^{c}}=\frac{i}{4}(\bar{\partial}-\partial)$ so that $\operatorname{\textrm{dd}^{c}}=\frac{i}{2}\partial\bar{\partial}$.
Let $L$ be a Hermitian holomorphic line bundle on $X$. The (singular) Hermitian metric on $L$ will be written as $h=e^{-\phi}$, and we will refer to such a symbol $\phi=-\operatorname{\textrm{dd}^{c}}\log h$ as a weight on $L$. In practice, $\phi$ is interpreted as a collection of locally integrable functions $\phi^{U}$ over trivializing open sets $U$ for $L$, called local weights. If $e^{U}$ is a local holomorphic trivializing section of $L$ over $U$ then $|e^{U}(z)|_{\phi}^{2}:=e^{-\phi^{U}(z)}$. We will use the weight $\phi$ instead of $h=e^{-\phi}$ for a Hermitian metric on $L$. The set of weights on $L$ is an affine space modeled on $L^{1}_{\rm loc}(X)$. Namely, if $\phi_{1}$, $\phi_{2}$ are two weights on $L$, their difference $\phi_{1}-\phi_{2}$ is a locally integrable function on $X$.
Equivalently, a weight can be seen as a function on the total space of the dual line bundle $L^{*}$ satisfying the log-homogeneity property $\phi(\lambda v)=\log|\lambda|+\phi(v)$ for all non-zero $v\in L^{*}$, $\lambda\in\mathbb{C}$. A section $s\in H^{0}(X,L)$ induces a weight on $L$ denoted by $\log|s|$, defined by $\log|s|(v):=\log|\langle s,v\rangle|$. The point-wise norm of $s$ in terms of the weight $\phi$ equals
$|s|^{2}_{\phi}:=|s|^{2}e^{-\phi}.$
If $\phi$ is in the class $\mathcal{C}^{2}$, i.e., it has continuous derivatives of order two the curvature form of the metric $\phi$ is the global form on $X$, locally expressed as $\operatorname{\textrm{dd}^{c}}\phi$ which represents the first Chern class $c_{1}(L)$ of $L$ in the second real de Rahm cohomology group of $X$. One should be careful with this suggestive notation since $\operatorname{\textrm{dd}^{c}}\phi$ is definitely not an exact form. The curvature form of a smooth metric is said to be positive if the local Hermitian matrix $(\phi_{i\bar{j}})=(\frac{\partial^{2}\phi}{\partial z_{i}\partial\bar{z}_{j}})$ is positive definite.
In general, the curvature $\operatorname{\textrm{dd}^{c}}\phi$ is still well-defined as a (1,1) current since the local weight $\phi^{U}$ is assumed to be locally integrable. The curvature current of a singular metric $e^{-\phi}$ is said to be positive if the local weights $\phi^{U}$ are plurisubharmonic (psh for short) functions. We interchangeably use $\omega_{\phi}$ and $\operatorname{\textrm{dd}^{c}}\phi$ for the curvature form/current of the weight $\phi$.
Given a positive measure $\mu$ on $X$ and a weight $\phi$ on $L$, we obtain $L^{2}$ norms on the space of holomorphic sections $H^{0}(X,L^{k})$ denoted by
$\|s\|^{2}_{k\phi,\mu}:=\int_{X}|s|^{2}e^{-k\phi}{\rm d}\mu.$
The Bergman kernel of order $k$ associated to $(L,\phi,\mu)$ is defined by for any $x\in X$
$\mathcal{K}_{k\phi,\mu}(x):=\sup\{|s(x)|^{2}:s\in H^{0}(X,L^{k}),\|s\|_{k\phi,\mu}\leq 1\}.$
Throughout the paper, we use $C$ for a positive constant whose value may change from line to line.
2.2. Bounded geometry on non-compact Kähler manifolds
We recall the definition of bounded geometry for Kähler manifolds (cf. [Kob84, CY80, WY20]). Denote by $B_{\mathbb{C}^{n}}(0,r)$ the open ball centered at the origin $0\in\mathbb{C}^{n}$ of radius $r>0$ with respect to the Euclidean metric $\omega_{\mathbb{C}^{n}}$.
Definition 2.1.
Suppose $(X,\omega)$ is a $n$-dimensional Kähler manifold. We say that $(X,\omega)$ has bounded geometry of order $k\in\mathbb{N}$ if there exists $0<r$ such that for any $p\in X$ there is a domain $U$ in $\mathbb{C}^{n}$ and a holomorphic map $\psi:U\rightarrow X$ of maximal rank everywhere, satisfying
  1. (1)
    $B_{\mathbb{C}^{n}}(0,r)\subset U$ and $\psi(0)=p;$
  2. (2)
    On $U$, there exists a constant $c>0$ depending only on $n$, $r$ such that
    $c^{-1}\omega_{\mathbb{C}^{n}}\leq\psi^{*}\omega\leq c\omega_{\mathbb{C}^{n}};$
  3. (3)
    there is a constant $A_{k}>0$ depending only on $k$ such that
    $\sup_{x\in B_{\mathbb{C}^{n}}(0,r)}\left|\frac{\partial^{|\alpha|+|\beta|}g_{i\bar{j}}(x)}{\partial z^{\alpha}\partial\bar{z}^{\beta}}\right|\leq A_{k},\;\;\forall\;|\alpha|+|\beta|\leq k,$
where $g_{i\bar{j}}$ is the component of $\psi^{*}\omega$ on $U$ in terms of natural coordinates $(z^{1},\ldots,z^{n})$, and $\alpha$, $\beta$ are the multiple indices with $|\alpha|=\alpha_{1}+\cdots+\alpha_{n}$.
The map $\psi$ is called a quasi-coordinate map and the pair $(U,\psi)$ is called a quasi-coordinate chart of $X$.
We have the following normal coordinates for a Kähler manifold of bounded geometry.
Lemma 2.2.
Suppose $(X,\omega)$ is a Kähler manifold with bounded geometry of order at least $5$. Then there exists $c>0$, $A_{k}>0$, for $k=1,2,\ldots$ such that for any point $p\in X$, there exist $0<\varepsilon=\varepsilon(p)<r$, and a holomorphic map $\kappa:B_{\mathbb{C}^{n}}(0,\varepsilon)\rightarrow X$ which is biholomorphic on its image with $\kappa(0)=p$, $\kappa^{*}\omega=dd^{c}\varphi$ satisfying the conditions (2), (3) in Definition 2.1, and $\varphi_{i\bar{j}}(0)=\delta_{ij}$, $\varphi_{i\bar{j}k}(0)=\varphi_{i\bar{j}\bar{k}}(0)=0$, and $\varphi_{i\bar{j}\bar{k}\bar{\ell}}(0)=\varphi_{ijk\bar{\ell}}(0)=0$.
Proof.
For any $p\in X$, we fix a local quasi-coordinate chart $(U,\psi)$ at $p$ as in Definition 2.1. By complex linear transformation we can assume that $\psi^{*}\omega=dd^{c}\phi$ on $U$ with $\phi_{i\bar{j}}(0)=\delta_{ij}$. Since $\psi$ is of maximal rank everywhere, there exists $0<\varepsilon_{1}=\varepsilon_{1}(p)<r$ such that $\psi:B(0,\varepsilon_{1})\rightarrow\psi(B(0,\varepsilon_{1}))\subset X$ is biholomorphic. We can choose $\varepsilon_{2}<\varepsilon_{1}$ depending on $A^{i}_{k\ell},B^{i}_{mnp}$ such that the holomorphic map $\tilde{\psi}:B(0,\varepsilon_{2})\rightarrow B(0,\varepsilon_{1})$, $z^{i}=\tilde{\psi}(w)^{i}=w^{i}+A^{i}_{k\ell}w^{k}w^{\ell}+B^{i}_{mnp}w^{m}w^{n}w^{p}$ is biholomorphic on its image, where the coefficients $A^{i}_{k\ell},B^{i}_{mnp}$ will be chosen hereafter.
Now we have $\kappa=\psi\circ\tilde{\psi}:B(0,\varepsilon_{2})\rightarrow X$ is biholomorphic on its image, and the pull-back metric $\kappa^{*}\omega=dd^{c}\varphi$ satisfies
$\varphi_{a\bar{b}}(w)$ $=\phi_{a\bar{b}}(\tilde{\psi}(w))+2A^{i}_{ak}\phi_{i\bar{b}}w^{k}+2\overline{A^{j}_{b\ell}\phi_{a\bar{j}}}\overline{w}^{\ell}+3B^{i}_{amn}w^{m}w^{n}\phi_{i\bar{b}}+3\overline{B^{j}_{bpq}}\phi_{a\bar{j}}\overline{w}^{p}\overline{w}^{q}$
$\quad+4A^{i}_{ak}\overline{A^{j}_{b\ell}}\phi_{i\bar{j}}w^{k}\overline{w}^{\ell}+O(|w|^{3}).$
Therefore we have
$\varphi_{a\bar{b}c}(0)=\phi_{a\bar{b}c}(0)+2A^{b}_{ac};\varphi_{a\bar{b}\bar{c}}(0)=\phi_{a\bar{b}\bar{c}}+2\overline{A^{b}_{ac}};\varphi_{a\bar{b}cd}(0)=\phi_{a\bar{b}cd}(0)+6B^{b}_{acd}+C,$
where $C$ only depends on $A^{i}_{jk}$ and $\phi_{a\bar{b},c}(0)$. By choosing $A^{b}_{ac}=-\frac{1}{2}\phi_{a\bar{b}c}(0)$ and then $B^{b}_{acd}=-\frac{1}{6}\phi_{a\bar{b}cd}(0)-C$, we obtain the local normal coordinates for any $p\in X$. In particular, $A^{a}_{bc}$ and $B^{a}_{bcd}$ are uniformly bounded for any $p\in X$ by the definition of bounded geometry. Since $\varepsilon_{2}<\varepsilon_{1}$ depends only on $A^{a}_{bc}$ and $B^{a}_{bcd}$, we get the map $\kappa:B(0,\varepsilon_{2})\rightarrow X$ as desired. ∎
3. Uniform convergence of Tsuji’s iteration
3.1. Asymptotic expansion of Bergman kernels
The asymptotic expansion of Bergman kernels (Theorem 3.1 below) in $C^{k}$-norms for Kähler manifolds with bounded geometry was established in Ma–Marinescu [MM07, Problem 6.1] (cf. also [MM15, Theorem 6]. In this section, we provide an alternative proof for the asymptotic expansion of Bergman kernels in $\mathcal{C}^{0}$-norm, which will be used later. Our proof is inspired by Donaldson [Don09] and Charles [Cha15] by using the Laplace approximation. We refer to [MM07, Section 6] for general results on the asymptotic expansion of the Bergman kernel on complete Kähler manifolds.
Theorem 3.1.
Let $(X,\omega_{\phi})$ be a complete Kähler manifold of dimension $n$, where $\omega_{\phi}:=dd^{c}\phi>0$ is the curvature of a Hermitian metric $h=e^{-\phi}$ on a line bundle $L$. Suppose that $(X,\omega_{\phi})$ has bounded geometry of order $\ell\geq 5$. Then for any $x\in X$,
$\left|B_{k\phi,\omega_{\phi}^{[n]}}-\left(\frac{k}{2\pi}\right)^{n}\left(1+\frac{b_{1}}{2k}\right)\right|_{\mathcal{C}^{0}}\leq Ck^{n-2}.$
where $\omega^{[n]}:=\omega^{n}/n!$, $b_{1}(x)$ is the scalar curvature of $\omega_{\phi}$, $C>0$ depending on the bounded geometry of $(X,\omega_{\phi})$, and $B_{k\phi,\omega_{\phi}^{[n]}}=\mathcal{K}_{k\phi,\omega_{\phi}^{[n]}}(x)e^{-k\phi(x)}$ denotes the Bergman function associated to $(k\phi,\omega_{\phi}^{[n]})$.
We will need the following Laplace approximation.
Lemma 3.2.
[Hör03, Theorem 7.7.5] Let $K\subset\mathbb{R}^{m}$ be a compact set, $U$ an open neighborhood of $K$ and $k$ a positive integer. If $\lambda>0$, $u\in C^{2k}(K)$ and $f\in C^{3k+1}(U)$, $f\leq 0$ in $U$, $f^{\prime}(x_{0})=0$, $f^{\prime}(x)\neq 0$ in $K\backslash\{x_{0}\}$, $\det H_{f}(x_{0})\neq 0$ then
[…]
for a constant $C$ depending on $\|f\|_{C^{3k+1}(U)}$, where $H_{f}$ denotes the Hessian matrix of $f$,
$L_{j}u=\sum_{s-r=j}\sum_{2s\geq 3r}2^{-s}\left\langle H_{f}(x_{0})^{-1}D,D\right\rangle^{s}(g^{r}_{x_{0}}u)(x_{0})/r!s!,$
with
$g_{x_{0}}(x)=f(x)-f(x_{0})-\frac{1}{2}\left\langle H_{f}(x_{0})(x-x_{0}),x-x_{0}\right\rangle.$
In particular, $L_{0}u=u(x_{0})$ and
$L_{1}u(x_{0})$ $=$ $\frac{1}{2}\Bigg{[}u\left\{-f_{ik\ell}f_{jrs}\left(\frac{1}{4}f^{ij}f^{k\ell}f^{rs}+\frac{1}{6}f^{ij}f^{ks}f^{r\ell}\right)+\frac{1}{4}f^{ij}f^{kl}f_{ijk\ell}\right\}$
$+f^{sq}f^{rp}f_{srq}u_{p}-{\rm Tr}(H_{u}H_{f}^{-1})\Bigg{]}_{x=x_{0}}.$
Proof of Theorem 3.1.
We will establish lower and upper bounds for the Bergman kernel by adapting the method of peak sections for compact Kähler manifolds, as developed in [Tia90] (see also [Rua98, Lu00, SW10]), and by using the Laplace method as in [Don09, Cha15].
For any $x_{0}\in X$, we take $(U,z=(z^{1},\ldots,z^{n}))$ a normal coordinate chart as in Lemma 2.2 centered at $x_{0}$. Then $L$ has a local holomorphic frame $g$ defined on $U$ on which the function $\psi=\phi-\log|g|^{2}=-\log|g|^{2}_{\phi}$ has minimum at $x_{0}$ and
(3.1) $\psi(z)=\sum_{j=1}^{n}|z^{j}|^{2}+\sum_{1\leq i,j,m,\ell\leq n}{\frac{1}{2!2!}}\phi_{i\bar{j}m\bar{\ell}}z^{i}\bar{z}^{j}z^{m}\bar{z}^{\ell}+O(|z|^{5}).$
Here for example, we can choose $a_{i},b_{ji},c_{\ell mi},d_{pqrs}$ such that $\log|g|^{2}(z)=\phi(0)+a_{i}z^{i}+b_{ji}z^{j}z^{i}+c_{\ell mi}z^{m}z^{\ell}z^{i}+d_{pqrs}z^{p}z^{q}z^{r}z^{s}$ and use Taylor’s expansion to have (3.1).
Let $\eta:[0,\infty)\to[0,1]$ be a cut-off function satisfying $\eta(t)=1$ for $t\leq 1/2$, $\eta(t)=0$ for $t\geq 1$, $|\eta^{\prime}(t)|\leq 4$ and $|\eta^{\prime\prime}(t)|\leq 8$. Define a $L^{k}$-valued $(1,0)$ form on $X$:
$\alpha=\bar{\partial}\left[\eta\left(\frac{k|z|^{2}}{(\log k)^{2}}\right)\right]g^{k},$
which vanishes outside $A_{k}:=\{z|(\log k)^{2}/(2k)\leq|z|^{2}\leq(\log k)^{2}/k\}\subset U$. Let $dV$ denote the Lebesgue measure on the local chart $(U,z)$. We define a weight function $\chi$, supported in $U$, by setting
$\chi=(n+2)\eta\left(\frac{k|z|^{2}}{2(\log k)^{2}}\right)\log\left(\frac{k|z|^{2}}{2(\log k)^{2}}\right)$
for $z\in U$, and $\chi=0$ outside $U$. We observe that $\chi\in\ [-2(n+2),0]$ on $A_{k}$ and can be approximated by a decreasing sequence of smooth functions $\{\chi_{m}\}_{m\in\mathbb{N}}$; cf. [Tia90, pages 104–105]. It follows from the bounded geometry property that for $k>0$ sufficiently large,
$\textrm{Ric}(\omega_{\phi})+k\omega_{\phi}+\operatorname{\textrm{dd}^{c}}\chi_{m}\geq\frac{k}{C}\omega_{\phi},$
for a uniform constant $C>0$. On the other hand, there exists a constant $C>0$ depending on $n$ and the bounded geometry of $(X,\omega_{\phi})$,
(3.2) $e^{-\chi}|\alpha|^{2}_{\omega_{\phi}}\leq C\left|\eta^{\prime}\left(\frac{k|z|^{2}}{(\log k)^{2}}\right)\right|^{2}\frac{k^{2}}{(\log k)^{4}}\phi^{i\bar{j}}z^{i}\bar{z}^{j}|g|^{2k}\leq C\frac{k|g|^{2k}}{(\log k)^{2}},$
on $A_{k}$, otherwise $\alpha=0$. Since on $A_{k}$, we have $e^{-k\psi}=(1-\sum_{j=1}^{n}|z^{j}|^{2})^{k}+O(|z|^{3})$ as follows from the choice of coordinates (3.1), hence locally
$e^{-k\psi}\omega_{\phi}^{n}\leq\left(1-\frac{1}{2}|z|^{2}\right)^{k}n!{\rm d}V,$
where ${\rm d}V=\left(\frac{i}{2}\right)^{n}dz^{1}\wedge d\bar{z}^{1}\wedge\cdots\wedge dz^{n}\wedge d\bar{z}^{n}$. It follows from (3.2) that
(3.3) $\int_{A_{k}}|\alpha|^{2}_{\omega_{\phi}}e^{-\chi-k\phi}\omega_{\phi}^{[n]}$ $\leq C\frac{k}{(\log k)^{2}}\int_{A_{k}}|g|^{2k}e^{-k\phi}\omega_{\phi}^{[n]}$
$=C\frac{k}{(\log k)^{2}}\int_{A_{k}}e^{-k\psi}\omega_{\phi}^{[n]}$
$\leq C\frac{k}{(\log k)^{2}}\left(\frac{(\log k)^{2}}{k}\right)^{n}\left(1-\frac{(\log k)^{2}}{k}\right)^{k}$
$\leq C(\log k)^{2n-2}k^{1-n-(\log k)/2}<\infty$
because $(1-(\log k)^{2}/k)k=e^{k\log(1-(\log k)^{2}/k)}\leq e^{-(\log k)^{2}/2}=k^{-(\log k)/2}$, where $C>0$ only depends on $n$ and the bounded geometry of $(X,\omega_{\phi})$.
We apply Hömander’s $L^{2}$ estimate (see [Hör73, Chapter 4] or [Tia90, Proposition 1.1]) with the weight $\chi$: there exists a smooth section $u$ of $L^{k}$ on $X$ such that ¯$\bar{\partial}u=\alpha$ and
$\int_{X}|u|^{2}e^{-k\phi-\chi}\omega_{\phi}^{[n]}$$\leq$ $\frac{C}{k}\int_{X}|\alpha|^{2}_{\omega_{\phi}}e^{-k\phi-\chi}\omega_{\phi}^{[n]}$
$=$ $\frac{C}{k}\int_{A_{k}}|\alpha|^{2}_{\omega_{\phi}}e^{-k\phi-\chi}\omega_{\phi}^{[n]}$
$\leq$ $C(\log k)^{2n-2}k^{-\log k/2-n},$
where the last inequality follows from (3.3). Because of the non-integrability of $e^{-\chi}$ at $x_{0}$, one must have $u(x_{0})=0$.
Since $\chi\leq 0$ we have
$\int_{X}|u|^{2}e^{-k\phi}\omega_{\phi}^{[n]}\leq\int_{X}|u|^{2}e^{-k\phi-\chi}\omega_{\phi}^{[n]}\leq C(\log k)^{2n-2}k^{-\log k/2-n}.$
It follows that $f:=\eta g^{k}-u$ is a holomorphic section of $L^{k}$, satisfying $|f(x_{0})|^{2}=e^{k\phi(0)}$ since $u(x_{0})=0$ and
$\|f\|^{2}_{k\phi}=\int_{X}|f|^{2}e^{-k\phi}\omega_{\phi}^{n}=\int_{U_{k}}|g|^{2k}e^{-k\phi}\omega_{\phi}^{n}+O(k^{-n-2}),$
where $U_{k}=\{z:|z|^{2}\leq(\log k)^{2}/k\}$. On $U_{k}$ we have
$\omega_{\phi}^{[n]}=(1+\phi_{i\bar{j}\ell\bar{\ell}}z^{i}\bar{z}^{j}+o(|z|^{3})){\rm d}V.$
Applying the Laplace approximation (Lemma 3.2) with $u=1+\phi_{i\bar{j}\ell\bar{\ell}}z^{i}\bar{z}^{j}+o(|z|^{3})$ and $f=-\psi$ where $\psi=-\log|g|^{2}_{\phi}$, we have
$\int_{U_{k}}|g|^{2k}e^{-k\phi}\omega_{\phi}^{[n]}$ $=$ $\int_{U_{k}}e^{-k\psi}\omega_{\phi}^{[n]}$
$=$ $\int_{U_{k}}e^{-k\psi}(1+\phi_{i\bar{j}\ell\bar{\ell}}z^{i}\bar{z}^{j}+o(|z|^{3})){\rm d}V$
$=$ $\left(\frac{2\pi}{k}\right)^{n}\left(1+\frac{1}{2k}\phi_{i\bar{i}j\bar{j}}+O(k^{-2})\right)$
$=$ $\left(\frac{2\pi}{k}\right)^{n}\left(1-\frac{1}{2k}S_{\omega_{\phi}}(0)+O(k^{-2})\right),$
where $O\left({k^{-2}}\right)$ dominated by $C/k^{-2}$ on $U$ with a uniform constant $C$ depending only on the bound of $|D^{\alpha}\phi|$ on $U$ with $|\alpha|\leq 7$. Therefore
$\mathcal{K}_{k\phi,\omega_{\phi}^{[n]}}(x_{0})\geq\frac{|f(x_{0})|^{2}}{\|f\|^{2}_{k\phi}}\geq\left(\frac{k}{2\pi}\right)^{n}e^{k\phi(0)}\left(1+\frac{1}{2k}S_{\omega_{\phi}}(0)+O(k^{-2})\right),$
hence we get a lower bound for $\mathcal{K}_{k\phi,\omega_{\phi}^{[n]}}e^{-k\phi}$.
We look for an upper bound for $\mathcal{K}_{k\phi,\omega_{\phi}^{[n]}}e^{-k\phi}$ following the strategy in [Cha15]. It suffices to work on the coordinate chart $(U,z)$ as above. Let $f$ be a local holomorphic section of $L^{k}$ on $U$. We claim that
$|f(0)/g^{k}(0)|^{2}\leq\frac{\int_{\Delta_{R}}|f/g^{k}|^{2}e^{-k\psi_{0}}\rho{\rm d}V}{\int_{\Delta_{R}}e^{-k\psi_{0}}\rho{\rm d}V},$
where $\Delta_{R}\subset U$ is a polydisc with radius $R$ and
$\psi_{0}=|z^{j}|^{2}+\phi_{i\bar{j}m\bar{\ell}}z^{i}\bar{z}^{j}z^{m}\bar{z}^{\ell},\;\rho=1+\phi_{i\bar{j}\ell\bar{\ell}}z^{i}\bar{z}^{j}.$
[…]
$\int_{\Delta_{R}}e^{-k\psi_{0}}z^{\alpha}\rho{\rm d}V=0.$
It follows that for any holomorphic function $h:U\to\mathbb{C}$,
$\int_{\Delta_{R}}|h|^{2}e^{-k\psi_{0}}\rho{\rm d}V=\int_{\Delta_{R}}\big{(}|h(0)|^{2}+|h-h(0)|^{2}\big{)}e^{-k\psi_{0}}\rho{\rm d}V.$
We thus apply with the holomorphic function $f/g^{k}$ (since $g$ has no zeros in the polydisc $\Delta_{R}$ of radius $R$ for $R>0$ sufficiently small) to obtain our claim.
Choosing $R=k^{-2/3}$, we have
$-k\psi_{0}\leq-k\psi+Ck^{-7/3}\;\text{and}\;\rho dV\leq(1+Ck^{-2})\omega_{\phi}^{[n]}$
hence
[…]
$=(1+Ck^{-2})\int_{\Delta_{R}}|f|^{2}e^{-k\phi}\omega_{\phi}^{[n]}.$
Using the Laplace approximation again, we have
[…]
Therefore we get
[…]
$=\left(\frac{k}{2\pi}\right)^{n}\left(1+\frac{1}{2k}S_{\omega_{\phi}}(0)+O(k^{-2})\right)\int_{\Delta_{R}}|f|^{2}e^{-k\phi}\omega_{\phi}^{[n]}.$
Since $|g^{k}(0)|^{2}=e^{k\phi(0)}$, we infer that
$|f(0)|^{2}\leq e^{k\phi(0)}\left(\frac{k}{2\pi}\right)^{n}\left(1+\frac{1}{2k}S_{\omega_{\phi}}(0)+O(k^{-2})\right)\int_{\Delta_{R}}|f|^{2}e^{-k\phi}\omega_{\phi}^{[n]},$
hence we get a upper bound for $\mathcal{K}_{k\phi,\omega_{\phi}^{[n]}}e^{-k\phi}$. ∎
3.2. Convergence of Tsuji’s iteration
Recall that $L\to X$ is a positively curved line bundle on $X$ equipped with a smooth Hermitian metric $\phi_{L}$. Let $\Omega$ be a smooth volume form on $X$. For any Hermitian metric $\phi$ on $L$, we denote by $\mu_{\phi}=\frac{1}{n!}e^{\phi-\phi_{L}}\Omega$, which is a well-defined positive measure on $X$. We prove the following lemma, which is a variant version of [Ber09b, Lemma 3].
Lemma 3.3.
Let $\phi_{\infty}$ be the solution of
$(dd^{c}\phi_{\infty})^{n}=e^{\phi_{\infty}-\phi_{L}}\Omega,\quad\omega_{\phi_{\infty}}=\operatorname{\textrm{dd}^{c}}\phi_{\infty}>0$
such that $(X,\omega_{\phi_{\infty}})$ has bounded geometry of order $\ell\geq 5$. Suppose that $C_{1}$, $C_{2}$ are two real numbers satisfying $C_{1}\leq\phi-\phi_{\infty}\leq C_{2}$. Then we have
(3.4) $\beta_{k}(\phi)-\phi_{\infty}\geq\frac{k-1}{k}C_{1}-\varepsilon_{k}$
and
(3.5) $\beta_{k}(\phi)-\phi_{\infty}\leq\frac{k-1}{k}C_{2}+\varepsilon_{k},$
where $\varepsilon_{k}=c/k^{2}$ with $c>0$ only depending on the bounded geometry of $\phi_{\infty}$. Moreover, if $\omega_{\phi_{\infty}}$ is a Kähler–Einstein metric with negative curvature, and $d_{k}$ is defined in (1.8), then we can obtain $\varepsilon_{k}=c/k^{3}$.
Proof.
Since $(dd^{c}\phi_{\infty})^{n}=e^{\phi_{\infty}-\phi_{L}}\Omega=n!\mu_{\phi_{\infty}}$, using the first order asymptotic of Bergman kernels for the weight $\phi_{\infty}$ (cf. Theorem 3.1), we have
$\mathcal{K}_{k\phi_{\infty},\mu_{\phi_{\infty}}}e^{-k\phi_{\infty}}$ $=$ $\mathcal{K}_{k\phi_{\infty},\omega_{\phi_{\infty}}^{[n]}}e^{-k\phi_{\infty}}$
$=$ $\left(\frac{k}{2\pi}\right)^{n}\left(1+O\left(\frac{1}{k}\right)\right).$
By definition $d_{k}=\left(\frac{k}{2\pi}\right)^{n}$ we have
(3.6) $\beta_{k}(\phi_{\infty})=\frac{1}{k}\log(\mathcal{K}_{k\phi_{\infty}}/d_{k})=\phi_{\infty}+\varepsilon_{k},$
where $\varepsilon_{k}=O(\frac{1}{k^{2}})$.
Similarly, if $\omega_{\infty}$ is a Kähler–Einstein metric with ${\rm Ric}(\omega_{\phi_{\infty}})=-\omega_{\phi_{\infty}}$, using the second-order asymptotic behavior of Bergman kernels for the weight $\phi_{\infty}$ (cf. Theorem 3.1) we get
$\mathcal{K}_{k\phi_{\infty},\mu_{\phi_{\infty}}}e^{-k\phi_{\infty}}$ $=$ $\mathcal{K}_{k\phi_{\infty},\omega_{\phi_{\infty}}^{[n]}}e^{-k\phi_{\infty}}$
$=$ $\left(\frac{k}{2\pi}\right)^{n}\left(1-\frac{n}{2k}+O\left(\frac{1}{k^{2}}\right)\right).$
Therefore, for $d_{k}=\left(\frac{k}{2\pi}\right)^{n}\max\left\{\frac{1}{2},(1-\frac{n}{2k})\right\}$ and $k>n$, we have
(3.7) $\beta_{k}(\phi_{\infty})=\frac{1}{k}\log(\mathcal{K}_{k\phi_{\infty}}/d_{k})=\phi_{\infty}+\varepsilon_{k},$
where $\varepsilon_{k}=O(\frac{1}{k^{3}}).$
Next, it follows from the fact that $\phi-\phi_{\infty}\geq C_{1}$ for some constant $C_{1}$, the extremal characterization of Bergman kernel implies that
(3.8) $\mathcal{K}_{k\phi}\geq e^{(k-1)C_{1}}\mathcal{K}_{k\phi_{\infty}}.$
Indeed, if we pick a section $0\neq\sigma\in H^{0}{(X,L^{k})}$ with $\int_{X}|\sigma|^{2}e^{-k\phi_{\infty}}\mu_{\phi_{\infty}}<+\infty$ then for any $x\in X$,
$\frac{|\sigma(x)|^{2}}{\int_{X}|\sigma|^{2}e^{-k\phi}{\rm d}\mu_{\phi}}\geq e^{(k-1)C_{1}}\frac{|\sigma(x)|^{2}}{\int_{X}|\sigma|^{2}e^{-k\phi_{\infty}}{\rm d}\mu_{\phi_{\infty}}}.$
Taking the supremum, we obtain the inequality (3.8). Therefore, combining with (3.6) (reps. (3.7)) we have
$\beta_{k}(\phi)-\phi_{\infty}\geq\frac{k-1}{k}C_{1}-\varepsilon_{k}$
for some positive $\varepsilon_{k}=O(\frac{1}{k^{2}})$ (reps. $\varepsilon_{k}=O(\frac{1}{k^{3}})$). Similarly, we have
$\beta_{k}(\phi)-\phi_{\infty}\leq\frac{k-1}{k}C_{2}+\varepsilon_{k}$
as required. ∎
Proof of Theorems 1.1 & 1.3.
We define by recurrence $\phi_{k+1}:=\beta_{k}(\phi_{k})$. For any $k\geq 1$, denote $C_{k}$ the best constant in the inequality
(3.9) $\phi_{k}-\phi_{\infty}\geq C_{k}.$
Lemma 3.3 thus implies that
$C_{k+1}\geq\frac{k-1}{k}C_{k}-\varepsilon_{k}.$
Since $\varepsilon_{k}$ is of order $1/k^{2}$, this implies that $kC_{k+1}$ is bounded from below uniformly by $O\big{(}\log k\big{)}$. By the same way, the last part of Lemma 3.3 provides a uniform lower bound for $k\widetilde{C}_{k+1}$ where $\widetilde{C}_{k}$ is the best constant such that
$\phi_{k}-\phi_{\infty}\leq\widetilde{C}_{k}.$
Since $C_{k}\leq\widetilde{C}_{k}$ by definition, we have $C_{k}=O\left(\frac{\log k}{k}\right)$ and $\widetilde{C}_{k}=O\left(\frac{\log k}{k}\right)$. So, we get the desired convergence with rate $\frac{\log k}{k}$.
By the same argument, we obtain the uniform convergence with rate $\frac{1}{k}$ when $\omega_{\phi_{\infty}}$ is a Kähler–Einstein metric. ∎
4. Variation of Kähler–Einstein metrics
4.1. Positivity of relative Bergman kernels
Setup 4.1.
Let $p:X\to Y$ be a holomorphic surjective map from a Kähler manifold $X$ of dimension $n+m$ to a complex manifold $Y$ of dimension $m$. Let $Y^{\circ}\subset Y$ denote the set of points that are not critical values of $p$ in $Y$ and set $X^{\circ}=p^{-1}(Y^{\circ})$.
If $p:X\to Y$ is in Setup 4.1, then the relative canonical bundle associated with $p$ is defined by $K_{X/Y}:=K_{X}-p^{*}(K_{Y})$. It follows from [Ber09a, BP08] that we have an isomorphism between $K_{X_{y}}$ and $(K_{X}-p^{*}K_{Y})|_{X_{y}}=K_{X/Y}|_{X_{y}}$. Let $t=(t_{1},\ldots,t_{m})$ be a local coordinates near a point in $Y$. Then these coordinates define a local $(m,0)$-form $dt=dt_{1}\wedge\ldots\wedge dt_{m}$, which trivializes $K_{Y}$. This trivialization induces a local section $\tilde{u}$ of $K_{X}$ over $X_{y}$ by $u=u\wedge p^{*}({\rm d}t)=u\wedge{\rm d}\tilde{t}$, where $\tilde{t}_{j}=t_{j}\circ p,\,j=1\ldots m$, are a functions on $X$. Conversely, given a local section $\tilde{u}$ of $K_{X}$, we can write $\tilde{u}=u\wedge{\rm d}\tilde{t}$ locally. Then, the restriction of $u$ to the fibers $X_{y}$ is uniquely defined as a local section of $K_{X_{y}}$. Thus, we call $u$ the trace of $\tilde{u}$ on $K_{X_{y}}$, denote by $\tau_{y}(\tilde{u})$. Then, the isomorphism is given by
(4.1) $\begin{split}K_{X_{y}}+p^{*}(K_{Y})|_{X_{y}}&\rightarrow K_{X}|_{X_{y}}\\ u\otimes p^{*}{\rm d}t&\mapsto u\wedge{\rm d}\tilde{t}\end{split}$
with the inverse map $\tilde{u}\mapsto\tau_{y}(u)\otimes p^{*}(dt)$. Hence the isomorphism $K_{X_{y}}\to K_{X/Y}|_{X_{y}}$ is given by $u\mapsto u\wedge p^{*}(dt)/dt$.
Let $\omega$ be a relative Kähler metric on $X$ such that $\omega_{y}:=\omega|_{X_{y}}$ is a Kähler metric on $X_{y}$ for $y\in Y^{\circ}$. Assume that $L$ is a line bundle on $X$ and $h_{L}=e^{-\phi_{L}}$ is a singular Hermitian metric on $L$ with positive curvature current. Let $E$ be the vector bundle (of infinite rank) over $Y$ with fiber $E_{y}:=H^{0}(X_{y},(K_{X_{y}}+L|_{X_{y}})\otimes\mathcal{I}(\phi_{L}|_{X_{y}}))$, for any $y\in Y^{\circ}$. Here, $\mathcal{I}(\phi)$ denotes the multiplier ideal sheaf of holomorphic functions on X that are square integrable against $e^{-\phi}$. Then the (singular) Hermitian metric $\phi_{L}$ on $L$ induces a Hilbert norm $\|\cdot\|_{y}$ on $E_{y}$ by
(4.2) $\|u\|^{2}_{y}:=\int_{X_{y}}c_{n}u\wedge\bar{u}e^{-\phi_{L}}=:\int_{X_{y}}|u|_{\omega}^{2}e^{-\phi_{L}}{\rm d}V_{X_{y},\omega}.$
Let $x\in X^{\circ}$, and $y=p(x)$. Let $z_{1},\ldots,z_{n+m}$ be local coordinates for a neighborhood $V\subset X^{\circ}$ of $x$ and $t_{1},\ldots,t_{m}$ local coordinates near $y$. For $u\in H^{0}(X_{y},K_{X_{y}}+L|_{X_{y}})$, we define $u^{\prime}$ by
$u^{\prime}{\rm d}z=u\wedge p^{*}{\rm d}t.$
The fiberwise Bergman kernel $\mathcal{K}(x)$ of $(K_{X/Y}+L)|_{X^{\circ}}$ with respect to this norm is defined by
(4.3) $\mathcal{K}(x)=\sup_{u\in E_{y},\|u\|^{2}_{y}\leq 1}|u^{\prime}(x)|^{2}.$
We remark that this concept of the Bergman kernel coincides with the classical one. Thus, $\phi_{X/Y}(x):=\log\mathcal{K}(x)$ provides a Hermitian metric $h=e^{-\phi_{X/Y}}$ on $K_{X/Y}$.
We establish the pseudo-effectivity criterion of the relative adjoint line bundle $K_{X/Y}+L$ by showing the log-plurisubharmonicity of the fiberwise Bergman kernel. The only assumptions we need to obtain a non-trivial positively curved metric are the positivity of $L$, together with the existence of an $L^{2}$ section of $K_{X/Y}+L$. The following theorem is a non-compact extension of [BP08, Theorem 0.1] (see also [PT18, Theorem 4.1.1], [ZZ18, Theorem 1.3] and [Cao17, Theorem 3.5.]).
Theorem 4.2.
Let $X$ be weakly pseudoconvex Kähler manifold of dimension $n+m$ and $Y$ a complex manifold of dimension $m$. Let $p:X\to Y$ be a surjective holomorphic map. Let $L$ be a line bundle on $X$ equipped with a singular Hermitian metric $h_{L}=e^{-\phi_{L}}$ satisfying
  • $dd^{c}\phi_{L}\geq 0$ in the sense of currents;
  • $H^{0}(X_{y_{0}},(K_{X_{y_{0}}}+L|_{X_{y_{0}}})\otimes\mathcal{I}(\phi_{L}|_{X_{y_{0}}}))\neq 0$ for some point $y_{0}\in Y^{\circ}$.
Denote $\mathcal{K}_{\phi_{L}}$ the relative Bergman kernel of the line bundle $(K_{X/Y}+L)|_{X^{\circ}}$ Then $\mathcal{K}_{\phi_{L}}$ has a semipositive curvature current and extends across $X\backslash X^{\circ}$ to a metric with semipositive curvature current on all of X.
A consequence of Theorem 4.2 is the positivity of the relative Bergman kernel.
Corollary 4.3.
Let $X$ be weakly pseudoconvex Kähler manifold and $Y$ a complex manifold. Let $p:X\to Y$ be a surjective holomorphic map. Assume that $H^{0}(X_{y_{0}},K_{X_{y_{0}}})\neq 0$ for some point $y_{0}\in Y^{\circ}$. Then, the relative pluri-canonical line bundle $K^{k}_{X/Y}$ admits a metric with a positive curvature current. Moreover, this metric equals the fiberwise $k$-Bergman kernel metric on the fiber of $p$.
Proof.
First, we take $L=\mathcal{O}_{X}$ with a constant metric. Then Theorem 4.2 implies that $\phi^{(1)}_{X/Y}:=\log\mathcal{K}$ defines a Hertmitian metric on $K_{X/Y}$ with semi-positive curvature. Using Theorem 4.2 recurrently with $L=K^{k}_{X/Y}$ with hermitian metric $\phi_{X/Y}^{(k)}$ for $k=1,2,\ldots$ we get a new metric $\phi_{X/Y}^{(k+1)}$ on $K_{X/Y}^{(k+1)}$ with positive curvature current defined by the natural logarithm of the associated Bergman kernel. ∎
The proof of Theorem 4.2 follows the same argument as in the compact case, cf. [BP08, PT18, ZZ18]. We provide a sketch of the proof here for the reader’s convenience. We need the following version of the Osawa–Takegoshi theorem, which can be deduced from [ZZ18, Theorem. 1.1] (see also [Cao17, BL16, GZ15b]).
Lemma 4.4.
Let $(X,\omega)$ be a weakly pseudoconvex Kähler manifold. Let $p:X\rightarrow\Delta_{r}=\{t:|t|<r\}\subset\mathbb{C}$ be a surjective holomorphic map. Suppose that $X_{0}=p^{-1}(0)$ is a submanifold of codimension $1$. Let $L$ be a holomorphic line bundle over $X$ equipped with a (possibly singular) Hermitian metric $\phi_{L}$ such that $\operatorname{\textrm{dd}^{c}}\phi_{L}\geq 0$. Let $u\in H^{0}(X_{0},K_{X_{0}}+L|_{X_{0}})$ such that $\int_{X_{0}}|u|^{2}_{\omega}e^{-\phi_{L}}{\rm d}V_{X_{0}}<\infty$. Then there exists $U_{r}\in H^{0}(p^{-1}(\Delta_{r}),K_{X/Y}+L)$ such that $U_{r}|_{X_{0}}=u\wedge p^{*}dt$ and
$\frac{1}{\pi r^{2}}\int_{p^{-1}({\Delta_{r})}}|U_{r}|^{2}_{\omega}e^{-\phi_{L}}{\rm d}V_{X,\omega}\leq\int_{X_{0}}|u|^{2}_{\omega}e^{-\phi_{L}}{\rm d}V_{X_{0},\omega}.$
Proof of Theorem 4.2.
We follow the same argument in the compact case (see [BP08, PT18, ZZ18]). Recall that for any $y\in Y^{\circ}$, setting $\phi_{y}=\phi_{L}|_{X_{y}}$, we define the $L^{2}$ norm on $u_{y}\in H^{0}(X_{y},K_{X_{y}}+L|_{X_{y}})$ by
$\|u_{y}\|^{2}:=\int_{X_{y}}|u_{y}|^{2}_{\omega}e^{-\phi_{y}}{\rm d}V_{X_{y},\omega}.$
Let $(t,z)=(t_{1},\ldots,t_{m},z_{1},\ldots,z_{n})$ be local coordinates near a point $x\in X$. Set $y=p(x)$. The Bergman kernel of $K_{X_{y}}+L|_{X_{y}}$ on $X_{y}$ ($y=y(t)$) is defined by
(4.4) $\mathcal{K}_{t}(z)=\sup\left\{|u_{y}^{\prime}(z)|^{2}:u_{y}\in H^{0}(X_{y},K_{X_{y}}+L|_{X_{y}}),\|u_{y}\|\leq 1\right\}$
where $u_{y}=u^{\prime}_{y}(z){\rm d}z\otimes e$ with $e$ a local frame of $L$ near $x$.
Remark that $p:X^{\circ}\to Y^{\circ}$ is a holomorphic submersion. To show that $\log\mathcal{K}_{t}(z)$ is psh with respect to $(t,z)$, one needs to prove the subharmonicity of $\log\mathcal{K}_{t}(z)$ along any complex line. Without loss of generality, we can assume that $Y^{\circ}=\Delta_{R}\subset\mathbb{C}$ is a disk ($m=1$).
The first property is the upper semi-continuity of $\log\mathcal{K}_{t}(z)$ with respect to $(t,z)$. Indeed, thanks to Demailly’s global regularization [Dem82, Section 9], we can find an increasing sequence of smooth metrics $(h_{\varepsilon}=e^{-\phi_{\varepsilon}})_{\varepsilon>0}$ on $L$, converging to the $h_{L}=e^{-\phi_{L}}$ in $L^{1}$ topology and satisfying $\operatorname{\textrm{dd}^{c}}\phi_{\varepsilon}\geq-\lambda_{\varepsilon}\omega$ with $\lambda_{\varepsilon}\searrow 0$ outside a pluri-polar set (the set of points at which the Lelong numbers are positive). Then we have the corresponding Bergman kernel $\mathcal{K}_{t,\varepsilon}$ is continuous. We proceed the same as in [PT18, Theorem 1.1] to obtain that $\log\mathcal{K}_{t,\varepsilon}(z)$ decreases to $\log\mathcal{K}_{t}(z)$ for $(t,z)\in U$ as $\varepsilon\searrow 0$, we are done.
We next show that $\log\mathcal{K}$ satisfies the sub-mean inequality on every holomorphic disc in $Y^{\circ}$, namely, for any $t_{0}\in Y^{\circ}$,
$\log\mathcal{K}_{t_{0}}(z)\leq\frac{1}{\pi r^{2}}\int_{\Delta_{r}}\log\mathcal{K}_{t}(z)i{\rm d}t\wedge{\rm d}\bar{t}$
for some $r<<1$ and $\Delta_{r}=\Delta(t_{0},r)\subset Y^{\circ}$. Fix $u_{0}\in H^{0}(X_{t_{0}},K_{X_{t_{0}}}+L|_{X_{t_{0}}})$ such that $\|u_{0}\|=1$ and $\mathcal{K}_{0}(z)=|u_{0}^{\prime}(z)|^{2}$. By the Ohsawa–Takegoshi type theorem (Lemma 4.4), $u_{0}$ can be extended to a holomorphic section $U$ on the preimage of $\Delta_{r}$ such that
$\frac{1}{\pi r^{2}}\int_{\Delta_{r}}\int_{X_{t}}|U^{\prime}(t,z)|^{2}e^{-\phi_{L}}{\rm d}V_{X_{t},\omega}i{\rm d}t\wedge{\rm d}\bar{t}\leq\|u_{0}\|^{2}=1,$
where we have used the Fubini theorem and $U=U^{\prime}dz\wedge p^{*}({\rm d}t)\otimes e$. Since $U(t,z)$ is holomorphic in $t$, we can write
$\log\mathcal{K}_{t_{0}}(z)=\log|U^{\prime}(0,z)|^{2}\leq\frac{1}{\pi r^{2}}\int_{\Delta_{r}}\log|U^{\prime}(t,z)|^{2}i{\rm d}t\wedge{\rm d}\bar{t}.$
Using the extremal property of the Bergman kernel, the term on the right-hand side is dominated by
$\frac{1}{\pi r^{2}}\int_{\Delta_{r}}\left(\log\mathcal{K}_{t}(z)+\log\int_{X_{t}}|U^{\prime}(t,z)|^{2}e^{-\phi_{L}}{\rm d}V_{X_{t},\omega}\right)i{\rm d}t\wedge{\rm d}\bar{t}.$
By Jensen’s inequality, we obtain
$\frac{1}{\pi r^{2}}\int_{\Delta_{r}}\log\int_{X_{t}}$ $|U^{\prime}(t,z)|^{2}e^{-\phi_{L}}{\rm d}V_{X_{t},\omega}i{\rm d}t\wedge{\rm d}\bar{t}$
$\leq\log\left(\frac{1}{\pi r^{2}}\int_{\Delta_{r}}\int_{X_{t}}|U^{\prime}(t,z)|^{2}e^{-\phi_{L}}{\rm d}V_{X_{t},\omega}i{\rm d}t\wedge d\bar{t}\right)\leq 0.$
This tells us that $\log\mathcal{K}_{t}(z)$ is psh in the horizontal direction. On the other hand, by the convexity of $|u_{t}^{\prime}(z)|$ and the definition of $\log\mathcal{K}$, we deduce that $\log\mathcal{K}(z)$ is also psh in the fiberwise direction. Therefore, $\log\mathcal{K}$ is psh on $p^{-1}(Y^{\circ})$, and in particular, the curvature of the induced singular Hermitian metric is semi-positive on $p^{-1}(Y^{\circ})$ in the sense of currents.
It remains to show that the curvature current of the induced Hermitian metric above can be extended over the whole of $X$. It suffices to show that its local potential is bounded from above near the singular fibers.
Fixing an open subset $V\subset X$, we define the function $\mathcal{K}$ on $V\cap X^{\circ}$ as (4.4). The problem is equivalent to proving that $\mathcal{K}$ is uniformly bounded from above on $V\cap X^{\circ}$. For any $x\in V\cap X^{\circ}$, there exists a section $\sigma\in H^{0}(X_{p(x)},K_{X_{p(x)}}+L|_{X_{p(x)}})$ such that
$\mathcal{K}(x)=|\sigma^{\prime}(z)|^{2},\quad\int_{X_{p(x)}}|\sigma|^{2}e^{-\phi_{p(x)}}{\rm d}V_{X_{p(x)},\omega}=1,$
where $(t,z)$ local coordinates near $x$ and $\sigma=\sigma^{\prime}dz\otimes e$. Using the Ohsawa–Takegoshi extension theorem, we can find an extension $\tilde{\sigma}$ of $\sigma$ on $p^{-1}(p(V))$ such that
$\int_{p^{-1}(p(V))}|\tilde{\sigma}|^{2}e^{-\phi_{L}}{\rm d}V_{X,\omega}\leq C_{V}$
with $C_{V}>0$ only depending on $V$. It follows from the sub-mean value inequality that $|\sigma^{\prime}(z)|^{2}$ is dominated by a constant only depending on $C_{V}$. This completes the proof. ∎
4.2. Plurisubharmonic variation of Kähler–Einstein metrics
This section is devoted to the study of the plurisubharmonic variation of Kähler–Einstein metrics on non-compact complex manifolds. The idea goes back to Tsuji’s work [Tsu10, Tsu13] in which he showed analogous results for families of canonically polarized compact complex manifolds and families of bounded pseudoconvex domains, using Berndtsson’s results [Ber09a, Ber06]. For alternative proofs in the compact case or on strong pseudoconvex domains, we refer the reader to [Sch12, Cho15b, Cho15a].
Theorem 4.5.
Let $p:X\to Y$ be in Setup 4.1. Assume further that $X$ is weakly pseudoconvex, and
  • (i)
    for every $y\in Y^{\circ}$, $X_{y}$ admits a Kähler–Einstein metric $\omega_{\textrm{KE},y}$ with negative Ricci curvature such that $(X_{y},\omega_{\textrm{KE},y})$ has bounded geometry of order $\ell\geq 5$ which is locally uniform in $y$, i.e. for any neighborhood $U\Subset Y^{\circ}$ of $y$ the constants $r,c,A_{k}$ in Definition 2.1 only depend on $U$.
  • (ii)
    $K_{X/Y}$ admits a singular Hermitian metric $h_{0}=e^{-\phi_{0}}$ with semi-positive curvature current and for every $y_{0}\in Y^{\circ}$, there exists neighborhood $U\Subset Y^{\circ}$ of $y_{0}$ such that $\|\phi_{0,y}-\phi_{{{\rm KE},y}}\|_{L^{\infty}(X_{y})}\leq C(U)$ for every $y\in U$.
Then the Hermitian metric on $(K_{X/Y})|_{X^{0}}$ induced by the Kähler–Einstein metrics on the fibers is positively curved. Moreover, if we have the additional condition that
  • (iii)
    the Kähler–Einstein potential $(\phi_{{\rm KE},y})_{y\in Y^{\circ}}$ is bounded from above near the singular fibers,
then this metric extends canonically across $X\setminus X^{\circ}$ to a positively curved metric.
The condition $(ii)$ can be satisfied if we can compare the Bergman kernel for $H^{0}(X_{y},K_{X_{y}})$ with the Kähler–Einstein measure $\omega_{{\rm KE},y}^{n}$. Indeed, in this case, the fiberwise Bergman kernel for $H^{0}(X_{y},K_{X_{y}})$ gives us a Hermitian metric $\phi_{0}$ on $K_{X/Y}$ which satisfies $(ii)$. We can see the following corollary as an example using the comparison of the Bergman kernel and the Kähler–Einstein measure in Greene–Wu [GW79].
Corollary 4.6.
Let $p:X\to Y$ be in Setup 4.1. Assume further that $X$ is weakly pseudoconvex and that
  • (i)
    for every $y_{0}\in Y^{\circ}$, there exist a neighborhood of $y_{0}$, $U\Subset Y^{\circ}$ and $0<A_{U}<B_{U}$ such that for any $y\in U$, $X_{y}$ admits a Kähler metric $\omega_{y}$ whose holomorphic sectional curvature belongs to $(-B_{U},-A_{U})$;
  • (ii)
    $X_{y}$ is simply-connected.
Then, the curvature current of the Hermitian metric on $(K_{X/Y})|_{X^{\circ}}$ induced by the Kähler–Einstein metrics on the fibers is positive.
Proof.
It follows from [WY20, Corollary 7] that for any $y\in U$, on $X_{y}$ admits the unique Kähler–Einstein metric which is uniformly equivalent to $\omega_{y}$ on $X$, where the equivalence constants depend only on $A_{U},B_{U}$ and ${\rm dim}X_{y}$. Moreover, by [GW79, p. 144, Theorem H (3)] the fiberwise Bergman kernel $\mathcal{K}_{y}$ for $H^{0}(X_{y},K_{X_{y}})$ satisfies $A_{1}\omega^{n}\leq\mathcal{K}_{y}\leq A_{2}\omega^{n}$ for some positive constants $A_{1},A_{2}$ depending only on $A_{U},B_{U}$ and ${\rm dim}X_{y}$. Therefore the fiberwise Bergman kernel for $H^{0}(X_{y},K_{X_{y}})$ provides us a Hermitian metric $\phi_{0}$ on $K_{X/Y}$ which satisfies $(ii)$ in Theorem 4.5. ∎
Proof of Theorem 4.5.
Fix $y\in Y^{\circ}$. We first observe that $H^{0}(X_{y},K^{(k_{0}+1)}_{X_{y}}\otimes\mathcal{I}(k_{0}\phi|_{X_{y}}))\neq 0$ for some $k_{0}>0$. Indeed, it follows from the Hömander $L^{2}$ estimate with the same argument as in the proof of Theorem 3.1, that there exists $k_{0}>0$ depending on the bounded geometry of $(X_{y},\omega_{{\rm KE},y})$ such that $H^{0}(X_{y},K^{(k_{0}+1)}_{X_{y}}\otimes\mathcal{I}(k_{0}\phi_{\rm KE}|_{X_{y}}))\neq 0$. By $(ii)$ we have $\|\phi_{0,y}-\phi_{{\rm KE},y}\|\leq C$, hence $H^{0}(X_{y},K^{(k_{0}+1)}_{X_{y}}\otimes\mathcal{I}(k_{0}\phi|_{X_{y}}))\neq 0$.
We now adapt the argument in [Tsu10]. Take $L_{k_{0}}:=K_{X/Y}^{k_{0}}$, then $h_{k_{0}}=e^{-\phi_{L_{k_{0}}}}$ with $\phi_{L_{k_{0}}}:=-k_{0}\phi_{0}$ is a singular Hermitian metric with with semi-positive curvature current. By iteration, at step $k\geq k_{0}$, assume that $L_{k}:=K_{X/Y}^{k}|_{X^{\circ}}$ admits a Hermitian metric $h_{k}=e^{-k\phi_{k}}$ with $\operatorname{\textrm{dd}^{c}}\phi_{k}\geq 0$, then we define $\phi_{k+1}:=\frac{1}{k}\log(\mathcal{K}_{k\phi_{k}}/d_{k})$ where $\mathcal{K}_{k\phi_{k}}$ is the Bergman kernel of $(K_{X/Y}+L_{k}{})|_{X^{\circ}}$. On each fiber $X_{y}$, we set $\phi_{k+1,y}=\beta_{k}(\phi_{k,y})=\frac{1}{k}(\log\mathcal{K}_{k\phi_{k,y}}/d_{k})$. By Theorem 3.1, we have $\|\phi_{k,y}-\phi_{{\rm KE},y}\|\leq C$, hence as above we have $H^{0}(X_{y},(K_{X_{y}}+L_{k}|_{X_{y}})\otimes\mathcal{I}(\phi_{L_{k}}|_{X_{y}}))\neq 0$. By Theorem 4.2 we infer that $\operatorname{\textrm{dd}^{c}}\phi_{k+1}\geq 0$.
Now, for any $y_{0}\in Y^{\circ}$, if we take a neighborhood $U\Subset Y^{\circ}$ of $y_{0}$ then it follows from the hypothesis and Theorem 1.3 that $\|\phi_{k,y^{\prime}}-\phi_{{\rm KE},y^{\prime}}\|_{L^{\infty}(X_{y^{\prime}})}\leq C(U)/k$ for all $y^{\prime}\in U$, hence $\|\phi_{k}-\phi_{\rm KE}\|_{L^{\infty}(p^{-1}(U))}\leq C(U)/k$. Therefore, the limit metric $\phi_{\rm KE}:=\lim_{k}\phi_{k}$ is a singular Hermitian metric on $K_{{X}/Y}$ with positive curvature current. ∎
4.3. Families of Kähler–Einstein metrics on bounded pseudoconvex domains
We recall some properties of pseudoconvex domains (cf. [CY80]). Let $\Omega$ be a strongly pseudoconvex domain in $\mathbb{C}^{n}$. Let $\rho$ be a smooth defining function for $\Omega$, i.e., $\rho\in\mathcal{C}^{\infty}(\overline{\Omega})$, $\Omega=\{\rho<0\}$, $\partial\rho\neq 0$ on $\partial\Omega$, and $(\rho_{i\bar{j}})>0$ in $\overline{\Omega}$. We define $\varphi_{\Omega}=-(n+1)\log(-\rho)$, which is a strictly plurisubharmonic function on $\Omega$ and $\varphi_{\Omega}(z)\rightarrow+\infty$ as $z\rightarrow\partial\Omega$. It follows from [CY80] that $(\Omega,\omega_{\varphi_{\Omega}})$ is a complete Kähler manifold, where $\omega_{\varphi_{\Omega}}=dd^{c}\varphi_{\Omega}$. Since
$\det((\varphi_{\Omega})_{i\bar{j}})=(n+1)^{n}\left(-\frac{1}{\rho}\right)^{n+1}\det(\rho_{i\bar{j}})(-\rho+|d\rho|^{2}),$
we have the Ricci curvature
${\rm Ric}(\omega_{\varphi_{\Omega}})=-(\varphi_{\Omega})_{i\bar{j}}+\partial_{i}\partial_{\bar{j}}F,$
where $F=-\log[\det(\rho_{\bar{j}i})(-\rho+|d\rho|^{2})]$ is smooth in $\Omega$ and bounded on $\overline{\Omega}$.
As explained in [CY80] that $(\Omega,\omega_{\varphi_{\Omega}})$ has bounded geometry. The following theorem is due to Cheng–Yau [CY80, MY83].
Theorem 4.7.
There exists a unique complete Kähler–Einstein metric $\omega_{\rm KE}=dd^{c}\phi_{\rm KE}$ with $-(n+1)$ scalar curvature which solves
(4.5) $(\operatorname{\textrm{dd}^{c}}\phi_{\rm KE})^{n}=e^{(n+1)(\phi_{\rm KE}-\phi_{\Omega})+F}\omega_{\varphi_{\Omega}}^{n},$
and $|\nabla^{k}_{\varphi_{\Omega}}(\phi_{\rm KE}-\varphi_{\Omega})|\leq C_{k}$ for any $k\in\mathbb{N}$. Moreover, $(\Omega,\omega_{\rm KE})$ has bounded geometry of any order.
Let $p:D\subset\mathbb{C}^{n}\times\mathbb{C}^{m}\to B\subset\mathbb{C}^{m}$ be the second projection where $D$ is a strictly pseudoconvex domain (with smooth boundary), $B$ is an open ball. Suppose that for each $t\in B$, $D_{t}:=p^{-1}(t)$ is bounded strongly pseudoconvex domain with smooth boundary. It follows from [CY80, MY83] (cf. Theorem 4.7) that on each slice $D_{t}$ there exists a unique complete Kähler form $\omega_{\phi_{t}}$ such that
$\operatorname{\textrm{Ric}}(\omega_{\phi_{t}})=-(n+1)\omega_{\phi_{t}}.$
Then we have the following corollary of Theorem 4.5:
Theorem 4.8.
The metric $\phi_{t}$ induces a Hermitian metric on the relative line bundle $K_{D/B}$ with positive curvature current
The result was proved by Choi [Cho15b, Cho15a] using the boundary behavior of the geodesic curvature, which satisfies a certain elliptic equation, and by Tsuji [Tsu13] using his construction of Kähler–Einstein metrics.
Proof.
Let $(z,t)\in\mathbb{C}^{n}\times\mathbb{C}^{m}$ be the standard coordinates. We denote by $\rho$ the defining function of $D$ which satisfies
  • $\rho\in\mathcal{C}^{\infty}(\overline{D})$ and $D=\{(z,t):\rho(z,t)<0\}$;
  • $\partial\rho\neq 0$ on $\partial D$,
  • $(\rho_{i\bar{j}}(\cdot,t))>0$ in $\overline{D}$.
This implies that $\varphi_{D}(\cdot,t)=-\log(-\rho(\cdot,t))$ is a strictly smooth plurisubharmonic function in $D_{t}$ for each $t\in B$. Applying Theorem 4.5, it suffices to show that $\|\phi_{t}-\varphi_{D_{t}}\|\leq C$ independent of $t$. Setting $u_{t}=\phi_{t}-\varphi_{D}(\cdot,t)$, it solves
$\det((\varphi_{D})_{i\bar{j}}(\cdot,t)+(u_{t})_{i\bar{j}})=e^{(n+1)u+F}\det((\varphi_{D})_{i\bar{j}}(\cdot,t))$
in $D_{t}$ where the function $F$ is bounded on $\overline{D}$ (defined as above). Applying Theorem 4.5, it suffices to show that $\|u_{t}\|_{L^{\infty}(D_{t})}\leq C$ uniformly in $t$. Indeed, it follows from the maximum principle [CY80, Proposition 4.1] that
$\sup_{D_{t}}|u_{t}|\leq\frac{1}{n+1}\sup_{D_{t}}|F|\leq C$
for a constant $C>0$ uniformly in $t$. This finishes the proof. ∎
4.4. Families of Kähler–Einstein metrics with cusp singularities
Let $p:\overline{X}\to Y$ be a holomorphic subjective proper map between compact Kähler manifolds with relative dimension $n$, $D=\sum_{i=1}^{\ell}D_{i}$ a reduced divisor with simple normal crossing support. Let $X:=\overline{X}\setminus D$. Assume $Y^{\circ}$ is the largest Zariki open subset of $Y$ such that if $\overline{X}^{\circ}:=p^{-1}(Y^{\circ})$ then every fiber ${\overline{X}}_{y}:=p^{-1}(\{y\})$ is smooth, $D_{y}:=D\cap{\overline{X}_{y}}$ has simple normal crossing support. We denote $X^{\circ}=\overline{X}^{\circ}\setminus D$.
Assume that $K_{\overline{X}_{y}}+D_{y}$ is ample for any $y\in Y^{\circ}$. By the well-known result of Kobayashi [Kob84] and Tian–Yau [TY87], building upon Cheng–Yau’s strategy [CY80], for each fiber, there exists a (unique) complete Kähler metric $\omega_{y}$ on $X_{y}:=\overline{X}_{y}\backslash D_{y}$ with cusp singularities along $D_{y}$ and $\operatorname{\textrm{Ric}}(\omega_{y})=-\omega_{y}$ on $X_{y}$. The metric $\omega_{y}$ can be extended as a positive current on $\overline{X}_{y}$; it satisfies
(4.6) $\operatorname{\textrm{Ric}}(\omega_{y})=-\omega_{y}+[D_{y}]$
in the sense of currents on $\overline{X}_{y}$, where $[D_{y}]$ is the integration current along $D_{y}$.
Then we will see in Theorem 4.11 below that a similar statement in Theorem 4.5 also holds in this setting, hence we get the following application:
Theorem 4.9.
Assume the setup above, then the fiberwise cusp Kähler–Einstein metrics $\omega_{y}$ satisfying (4.6) can be glued to define a current $T$ on $\overline{X}$ such that:
  1. (i)
    $T$ is positive on $X^{\circ}=\overline{X}^{\circ}\backslash D$ ;
  2. (ii)
    $T$ extends canonically across $(X\backslash X^{\circ})\cap D$ to a closed positive current.
Proof.
If one chooses a background Kähler metric $\omega$ on $\overline{X}$ then it induces a Hermitian metric $h^{\omega}_{\overline{X}/Y}=e^{-\phi_{\omega}}$ on $K_{\overline{X}/Y}$ with curvature form $\Theta_{h^{\omega}}(K_{X_{y}})=dd^{c}\phi_{\omega}$, cf. [Pău17, §3.1]. Fix smooth Hermitian metrics $h_{i}=h_{D_{i}}=e^{-\phi_{i}}$ on $\mathcal{O}_{\overline{X}}(D_{i})$ with curvature form $\theta_{D_{i}}=dd^{c}\phi_{i}\in c_{1}(D_{i})$. Denote $\theta_{i,y}=\sum\theta_{D_{i}|_{\overline{X}_{y}}}$ for every generic point $y\in Y^{\circ}$. Let $s_{i}$ be a holomorphic section cutting out $D_{i}$, and denote $|s_{i}|^{2}=|s_{i}|^{2}_{h_{i}}$ and $|s|^{2}=\prod|s_{i}|^{2}$. On each fiber $\overline{X}_{y}$, we set $\eta_{y}=\Theta_{h^{\omega}}(K_{X_{y}})+\sum\theta_{D_{i}|_{X_{y}}}$ being chosen to be Kähler (the restriction of $\Theta_{h^{\omega}}(K_{\overline{X}/Y})+\sum\theta_{D_{i}}$). The equation (4.6) can be written as
(4.7) $(\eta_{y}+\operatorname{\textrm{dd}^{c}}\varphi_{y})^{n}=|s|^{-2}{e^{\varphi_{y}}\omega^{n}},$
where $\varphi_{y}$ is the unknown function. By [BG14, Thm. 4.2], there exists a unique $\varphi_{y}\in\mathcal{E}^{1}(\overline{X}_{y},\eta_{y})$, the space of functions with finite energy. We can glue the functions $\varphi_{y}$ to obtain a function $\varphi$ on $\overline{X}$. Thus, we obtain a closed $(1,1)$ current on $\overline{X}$
$T:=\Theta_{h^{\omega}}(K_{\overline{X}/Y})+\sum\theta_{D_{i}}+\operatorname{\textrm{dd}^{c}}\varphi=dd^{c}\phi_{KE}\in c_{1}(K_{\overline{X}/Y}+D),$
where $\phi_{\rm KE}:=\phi_{\omega}+\sum_{i}\phi_{i}+\varphi$. We will show that $T$ is positive on $X^{\circ}$ and extends across $X\backslash X^{\circ}$ to a closed positive current by applying Theorem 4.5.
It is convenient to work with the complete Kähler metric with Poincaré singularities $\omega_{P,y}=\eta_{y}-\sum_{i}\operatorname{\textrm{dd}^{c}}\log(\log|s_{i}|^{2})^{2}$ on $X_{y}=\overline{X}_{y}\backslash D_{y}$ (up to scaling the hermitian metrics on $\mathcal{O}_{\overline{X}}(D_{i})$). It indeed defines a smooth Kähler metric with Poincaré growth along $D_{y}$, having bounded geometry at any order on $X_{y}=\overline{X}_{y}\backslash D_{y}$ as follows from [Gue14, §1] (cf. also [Kob84, Lemma 2]); all of these properties are satisfied uniformly in $y$. Setting $\psi_{y}=\varphi_{y}+\sum_{i}\log(\log|s_{i}|^{2})^{2}$, we have
$(\omega_{{\rm P},y}+\operatorname{\textrm{dd}^{c}}\psi_{y})^{n}=e^{\psi_{y}+F}\omega_{{\rm P},y}^{n}$
where $F=\log\left(\frac{\omega^{n}}{\prod_{i}|s_{i}|^{2}\log|s_{i}|^{2}\omega_{\rm P}^{n}}\right)$ is known to be a bounded smooth function on $\overline{X}\backslash D$ (which is even smooth in the quasi-coordinates); cf. [Gue14, Lemma 1.6] or [CG72, Proposition 2.1].
For any $y_{0}\in Y^{\circ}$, take a neighborhood $U\Subset Y^{\circ}$ of $y_{0}$. By the result of Kobayashi [Kob84] and Tian–Yau [TY87], we have that $\|\psi_{y}\|_{L^{\infty}(X_{y})}\leq C(U)$ for all $y\in U$. Indeed, since the bisectional curvature of $(X_{y},\omega_{P,y})$ is bounded below by some constant depending on $U$ for all $y\in U$ (cf. [Gue14]), the Yau generalized maximum principle (see [CY80, Prop. 4.1]) yields
$\sup_{X_{y}}|\psi_{y}|\leq\sup_{X_{y}}|F|\leq C(U)$
for any $y\in U$ because $F$ is bounded on $\overline{X}\backslash D$. Hence, for $\phi_{0}:=\phi_{\omega}+\sum_{i}\phi_{i}-\sum_{i}\log(\log|s_{i}|^{2})^{2}$ and $\phi_{{\rm KE}}=\phi_{\omega}+\sum_{i}\phi_{i}+\varphi$, we have
$\sup_{X_{y}}|\phi_{{\rm KE},y}-\phi_{0,y}|=\sup_{X_{y}}|\psi_{y}|\leq C(U).$
Also, the result of [Kob84, TY87] shows that the Kähler–Einstein metric $dd^{c}\phi_{{\rm KE},y}=\eta_{y}+\operatorname{\textrm{dd}^{c}}\varphi_{y}$ is complete and has bounded geometry on $X_{y}=\overline{X}_{y}\backslash D_{y}$ which is locally uniform in $y\in Y^{\circ}$.
In our setting, we apply Theorem 4.11 with $\phi_{0}=\phi_{\omega}+\sum_{i}\phi_{i}-\sum_{i}\log(\log|s_{i}|^{2})^{2}$ and $\phi_{\rm KE}$, to infer that $T$ is a positive current on $X^{\circ}$. Moreover, we also have that $\phi_{{\rm KE}}$ is bounded from above near the singular fibers of $p|_{X}$. This can be proved following the same argument as in [Gue20, Section 6], which is based on the approach of Păun [Pău17, §3.3]. Therefore, we can extend $T$ to $X$ as a positive current.
We conclude the proof by showing that $\phi=\phi_{\rm KE}$ extends across $D_{\rm reg}$, and hence across $D$, following  [BG14, Lemma 3.7]. Fix $x\in D_{\rm reg}$, and let $(U,(z^{0},\ldots,z^{n}))$ be local coordinates around $x$ in $\mathbb{C}^{n+1}$. In these coordinates, we write $D=(z^{1}\cdots z^{\ell}=0)$. It suffices to treat the case $\ell=1$, as the general case can be proven similarly. We have the integral
$\int_{U\backslash D}\frac{e^{\phi}{\rm d}V}{|z^{1}|^{2}}\leq C\int_{0<|z^{1}|<1}\frac{i{\rm d}z^{1}\wedge{\rm d}\bar{z}^{1}}{|z^{1}|^{2}\log^{2}(|z^{1}|^{2})}<\infty.$
Suppose, by contradiction, that $\phi$ does not tend to $\infty$- ∞ near $D$. Then, there exists a sequence $(z_{k})$ in $U\setminus D$ converging to $D$ such that $\phi(z_{k})\geq-C$ for some constant $C$. Let $\Delta_{k}$ denote the polydisk centered at $z_{k}$ with polyradius $(r_{k},\delta,\ldots,\delta)$ for $r_{k}\searrow 0$ and some fixed $\delta>0$, so that $\Delta_{k}\subset U\setminus D$. It follows from the sub-mean value inequality for $\phi$ at $z_{k}$ that
$-C\leq\frac{1}{\operatorname{{\rm Vol}}(\Delta_{k})}\int_{\Delta_{k}}\phi{\rm d}V.$
Using Jensen’s inequality, we obtain
$e^{-C(n)}\leq\int_{\Delta_{k}}\frac{e^{\phi}{\rm d}V}{r_{k}^{2}\delta^{2n}}$
and on $\Delta_{k}$, $|z^{1}|\leq 2r_{k}$ hence
$C^{\prime}\leq\int_{\Delta_{k}}\frac{e^{\phi}{\rm d}V}{|z^{1}|^{2}}$
with $C^{\prime}=\frac{1}{4}e^{-C}\delta^{2n}$. Since the measure of $\Delta_{k}$ goes to zero as $k\to+\infty$ we infer that $\int_{U\backslash D}\frac{e^{\phi}dV}{|z^{1}|^{2}}$ is infinite, which is a contradiction. ∎
To end the section, we establish an extension result of the Ohsawa-Takegoshi type for the family of quasi-projective manifolds, which ensures that the statement in Theorem 4.2 and 4.5 holds in this case.
Lemma 4.10.
Let $(\overline{X},\omega)$ be a Kähler manifold fibered over the disc $\Delta_{r}=\{t\in\mathbb{C}:|t|<r\}$ with $r>1$, with compact fibers $\overline{X}_{t}$. Let $D$ be a reduced divisor with simple normal crossing support in $\overline{X}$ not contained in $\overline{X}_{0}$. Let $X=\overline{X}\backslash D$ and $X_{t}=\overline{X}_{t}\backslash D$. Let $L$ be a holomorphic line bundle on $\overline{X}$ equipped with a singular Hermitian metric $e^{-\phi}$ with semi-positive curvature. Let $u\in H^{0}(X_{0},K_{X_{0}}+L|_{X_{0}})$ such that $\int_{X_{0}}u\wedge\bar{u}e^{-\phi}<\infty$. Then there exists $U\in H^{0}(X,K_{X}+L|_{X})$ such that $U=u\wedge dt$ for $t=0$ and
(4.8) $\frac{1}{\pi r^{2}}\int_{X}U\wedge\bar{U}e^{-\phi}\leq\int_{X_{0}}u\wedge\bar{u}e^{-\phi}.$
Proof.
The proof is almost identical to that of [Cao17, Theorem 1.3], so we will be somewhat brief and focus on the differences. We observe that $X=\overline{X}\backslash D$ admits a complete metric $\omega_{\rm P}$ with Poincaré singularity type; cf. [Kob84]. We take a covering of $\overline{X}$ by finite open sets $V_{\alpha}$ biholomorphic to the polydisc in $\mathbb{C}^{n}$. For each $\alpha$ we fix a local coordinates $(t,z^{1}_{\alpha},\ldots,z^{n}_{\alpha})$ such that $\overline{X}_{0}\cap V_{\alpha}=(t=0)$ and $D\cap V_{\alpha}=(z_{\alpha}^{1}\cdots z_{\alpha}^{\ell}=0)$. So that $X\cap V_{\alpha}\simeq(\Delta^{*})^{\ell}\times\Delta^{n+1-\ell}$ and $X_{0}\cap V_{\alpha}\simeq(\Delta^{*})^{\ell}\times\Delta^{n-\ell}$ where $\Delta$ is the unit disc in $\mathbb{C}$ and $\Delta^{*}=\Delta\backslash\{0\}$. Locally on $U_{\alpha}$, we can write $u=f_{\alpha}\otimes e_{\alpha}$ where $f_{\alpha}$ is a holomorphic function on $X_{0}\cap V_{\alpha}$. We write the Hermitian metric $h=e^{-\phi}=h_{0}e^{-\varphi}$ for $h_{0}$ a fixed smooth Hemitian metric and a function $\varphi$ on $\overline{X}$. Thanks to Demailly’s approximation theorem, there exists an increasing sequence $\{h_{k}\}=\{h_{0}e^{-\varphi_{k}}\}$ of metrics with analytic singularities with curvature $i\Theta_{h_{k}}(L)\geq-\frac{\omega}{k}$.
Step 1. We claim that there exists a smooth section $\tilde{u}$ on $K_{X}+L|_{X}$ such that $\tilde{u}|_{X_{0}}=u$, ¯$\bar{\partial}\tilde{u}|_{X_{0}}=0$ and for some $\sigma>0$,
$\int_{X}\frac{|\bar{\partial}\tilde{u}|^{2}_{\omega_{\rm P},h_{0}}e^{-(1+\sigma)\varphi}}{|t|^{2}(\log|t|)^{2}}{\rm d}V_{X,\omega_{\rm P}}\leq C\int_{X_{0}}|u|^{2}_{\omega_{\rm P},h}{\rm d}V_{X_{0},\omega_{\rm P}}.$
Thanks to the strong openness [GZ15a], we can find $\sigma>0$ such that
$\int_{X_{0}\cap V_{\alpha}}|u|^{2}_{\omega_{p},h_{0}}e^{-(1+\sigma)\varphi}{\rm d}V_{X_{0},\omega_{\rm P}}\leq 2\int_{X_{0}\cap V_{\alpha}}|u|^{2}_{\omega_{\rm P},h}{\rm d}V_{X_{0},\omega_{\rm P}}$
We apply the usual version of the Ohsawa–Takegoshi theorem [Dem12, Theorem 13.6] with the weight $e^{-(1+\sigma)\varphi}$, we can find holomorphic sections $\tilde{f}_{\alpha}$ of $K_{X}+L|_{X}$ on $U_{\alpha}$ extending $u$ such that
$\int_{V_{\alpha}\backslash D}\frac{|\tilde{f}_{\alpha}|^{2}_{\omega_{\rm P},h_{0}}e^{-(1+\sigma)\varphi}}{|t|^{2}(\log|t|)^{2}}{\rm d}V_{X,\omega_{\rm P}}\leq\int_{X_{0}\cap V_{\alpha}}|u|^{2}_{\omega_{\rm P},h}{\rm d}V_{X_{0},\omega_{\rm P}}.$
Let $\{\chi_{\alpha}\}$ be a partition of unity subordinate to $\{V_{\alpha}\}$. We can check $\tilde{u}:=\sum_{\alpha}\chi_{\alpha}\tilde{f}_{\alpha}$ satisfies the claim.
Step 2. Set $\Psi=\log|t|^{2}$ and $g_{m}=\bar{\partial}((1-b_{m}^{\prime}\circ\Psi)\tilde{u})$ where $b_{m}$ is a $\mathcal{C}^{1}$ function defined on $\mathbb{R}$ such that
$b_{m}(s)=s\;\text{for}\;s\geq-m,\quad b_{m}^{\prime\prime}(s)=\mathbf{1}_{\{-m-1\leq s\leq-m\}}(s).$
We claim that there exists a sequence $\{a_{m}\}\subset\mathbb{N}$ tending to $\infty$, $L$-valued locally $L^{2}$-forms $\gamma_{m}$ and $\beta_{m}$ such that
¯$\bar{\partial}\gamma_{m}+(\frac{m}{a_{m}})^{\frac{1}{2}}\beta_{m}=g_{m},\quad\lim_{m\to\infty}\frac{m}{a_{m}}=0$
and for $k\geq a_{m}$
[…]
The proof of the claim follows line by line from the one in [Cao17, p. 25]. We note here that if $\varphi_{k}$ has singularities along the analytic set $Z_{k}$ then $X\backslash Z_{k}=\overline{X}\backslash(D\cup Z_{k})$ has a complete Kähler metric (cf. [Dem12, Lemma 12.9]), so we can apply the $L^{2}$ existence theorem for solutions to $\bar{\partial}$-equation; cf. [Dem12, Proposition 13.4, Remark 13.5], also [Cao17, Appendix]. While, in [Cao17], $X$ is assumed to be weakly pseudoconvex, so that $X\backslash Z$ is a complete Kähler manifold for an analytic subset $Z$. The latter property is still true in our setting.
Step 3. Up to passing to a subsequence, we assume that $((1-b_{m}^{\prime}\circ\Psi)\tilde{u}-\gamma_{m})_{m}$ converges weakly to some $U\in L^{2}(X,K_{X}+L|_{X})_{\omega_{p},h_{0}}$. It suffices to verify that $U$ is holomorphic and satisfies (4.8). Since ¯$\bar{\partial}((1-b_{m}^{\prime}\circ\Psi)\tilde{u}-\gamma_{m})=(m/a_{m})^{\frac{1}{2}}\beta_{m}$ converges to zero and $\bar{\partial}$ is a closed operator it follows that $F$ is holomorphic. The inequality (4.8) follows immediately from the one (4.9) (cf. [Cao17, page 26]).
It remains to check that $F|_{X_{0}\cap V_{\alpha}}=f$. We solve the $\bar{\partial}$-equation on $V_{\alpha}\backslash D$: ¯$\bar{\partial}w_{m}=\beta_{m}$ and
$\int_{V_{\alpha}\backslash D}|w_{m}|^{2}_{\omega_{\rm P},h_{a_{m}}}{\rm d}V_{X,\omega_{\rm P}}\leq C\int_{V_{\alpha}\backslash D}|\beta_{m}|^{2}_{\omega_{\rm P},h_{a_{m}}}{\rm d}V_{X,\omega_{\rm P}}\leq C_{1}$
for some uniform constant $C_{1}>0$. Then
$U_{m}:=(1-b_{m}^{\prime}\circ\Psi)\tilde{u}-\gamma_{m}-(\frac{m}{a_{m}})^{\frac{1}{2}}w_{m}$
is a holomorphic function on $V_{\alpha}\backslash D$, and $U_{m}\to U$ weakly. Since $U_{m}|_{X_{0}\cap V_{\alpha}}=u$ we have $U|_{X_{0}\cap V_{\alpha}}=u$. ∎
Theorem 4.11.
Theorems 4.2 and 4.5 hold in the case when $X$ is a complex manifold fibered over a complex manifold $Y$, with fiber $X_{y}=\overline{X_{y}}\backslash D$.
Proof.
The statement in Theorem 4.2 for the family $p:X=\overline{X}\setminus D\rightarrow Y$ follows immediately from the global regularization result and the Ohsawa-Takegoshi theorem above.
We argue the same as in Theorem 4.5, noticing that in Tsuji’s iteration process, the metric $\phi_{k}$ on $K_{X/Y}$ is bounded from above (which is equivalent with the metric $\phi_{\rm P}$ of Poincaré type). ∎